A Ferris wheel with a radius of 8.0 m makes 1 revolution in 10 s. When a passenger is at the top, he releases a ball. How far from the point on the ground directly under the release point does the ball land?
1
Expert's answer
2020-11-06T10:40:52-0500

The beginning O of plane coordinate system "{O}_{xy}"

 is located on the ground under the point of release. 


"O_y" axis is directed upward, "O_ x" axis has horizontal direction in the plane of the Ferris wheel. 


In such coordinate system ball has coordinates:


"x(t)=\u03c9rt~~~~~~~-(1)"

"y(t)=2R−\dfrac{gt^2}{2}~~~~-(2)"



here "ω=\dfrac{2\pi}{T}" and "T=10s=0.63rads^{-1}"


R is the radius of the Ferris wheel, "g=9.81m\/s^2"

 

 

Let t –time after release. If the ball fallen, 

then y = 0. 

also 

"t=2\sqrt{\dfrac{R}{g}}~~~~~~~-(3)"

Putting 3 in 1 we get,


"x=2ωR\sqrt{\dfrac{R}{g}}=2\times 0.63\times 8\sqrt{\dfrac{8}{9.81}}=9.1m"